# The SRG/eROSITA All-Sky Survey

### v1

@@
N. Zimmermann 11    M. Kluge 22    S. Grandis 11    T. Schrabback 11    F. Balzer 22    E. Bulbul 22    J. Comparat 22    B. Csizi 11    V. Ghirardini 22    H. Jansen 11    F. Kleinebreil 11    A. Liu 22    A. Merloni 22    M. E. Ramos-Ceja 22    J. Sanders 22    X. Zhang 22    P. Aschenbrenner 11    F. Enescu 11    S. Keiler 11    M. Märk 11    M. Rinner 11    P. Schweitzer 11    E. Silvestre-Rosello 11    L. Stepman 11
@@ 3.1 Radial colour distribution
where $\mathcal{L_{\mathrm{tot,none}}}$ is the total log-likelihood assuming no additional cluster component. Here, we follow the formalism for a one dimensional Gaussian likelihood, where $\ln{\mathcal{L_{\mathrm{Gauss,1D}}}}=-\frac{1}{2}\left(\frac{\Delta x}{\sigma}\right)^{2}$ and $\ln{\mathcal{L_{\mathrm{Gauss,1D}}^{\mathrm{ref}}}}=\ln{\mathcal{L_{\mathrm{Gauss,1D}}}}\left(\Delta x=0\right)$. Thus we define this quantity in terms of $\sigma$. Only sources with a detection significance $\mathrm{SIGN}\geq 3.0\sigma$ are treated as significant detections. In the case of a clear peak in the posterior distribution, we estimate the uncertainty for each parameter by taking the symmetric $68\%$ interval in parameter space around the retrieved maximum log-likelihood value. If $\mathrm{SIGN}<3\sigma$, we take the $95\%$-percentile of the retrieved posterior distribution as an upper limit on the respective parameter. If the peak of the posterior distribution is within $1\sigma$ of the prior boundary, we report the $5\%$-percentile as a lower boundary (if within $1\sigma$ of the upper prior boundary), and the $95\%$-percentile as an upper boundary (if within $1\sigma$ of the lower prior boundary). We validate the inference framework on mock catalogues, discussed in more detail in Appendix A
@@ 4 Results
In this section, we present the results of the method presented in Sect. 3.1 and the photometric redshift estimation described in Sect. 3.2. Additionally, we provide comments on the visual inspection of the targets analysed in this work. We show a summary of the significances found using the different colour indicies in Fig. 4, using both the X-ray centres and the optical centres provided by Kluge et al. (2024). In addition, Table 3 provides detailed results for the optical centres and iJi-J colours. The findings for the other combinations of cluster centre and colour are provided in Appendix C. Listed in the tables are the maximum posterior values or the upper/lower limit values ($95\%$ / $5\%$-percentile of the posterior distribution) of each of the parameters of our model, the significance of our detection as described in Sect. 3.1, a photometric redshift estimate obtained in this work ($z_{\mathrm{phot},i-J}$, iJi-J indicating the used colour index), the number of sources found to have a cluster member probability higher than $80\%$ ($n_{\mathrm{clm}}^{0.8}$), as well as the best redshift estimate computed by Kluge et al. (2024).
@@ 4.1 Radial colour over-density
We compare the estimated cluster member probabilities from the CAHA data and the the WST data in Fig. 8. The most probable cluster members (upper right corner of Fig. 8) are identified with both of the data sets, whereas the estimated cluster member probability of less probable cluster member galaxies scatters symmetrically around the one-to-one relation. For an image comparison, please see Fig 12 in Appendix B.
@@ 4.3 Visual inspection
For illustration we provide image cutouts around the respective centres (see Appendix C), and comment on each cluster candidate, as well as our respective findings. We order the targets in three different classes defined by a visual assessment and their statistical significance emerging from our analysis:
@@ 5.3 Cluster centre proxies
We test the eRASS1 X-ray centres as reported in Bulbul et al. (2024) (see Appendix C), as well as optical centres found in Kluge et al. (2024) (see Table 3 and Appendix C). In principle, the optical centre is a more suitable centre proxy for our purposes because most of the times it is located at the position of a high probability, bright cluster member galaxy found in earlier works (Kluge et al. 2024). The colour of the central galaxy has an influence of the extracted mean colour, due to the designed radial weighting of the sources’ colours with the weighting function $f(A,R_{i})$ in Eq. (2). Thus, if the cluster centre is placed on an actual cluster member galaxy, the measured colour is weighted accurately by $f(A,R_{i})$ to find the correct red-sequence colour. If the cluster centre is wrongly placed on a random field galaxy, the measured colour does not correspond to the correct red-sequence colour and is mistakenly weighted strongly by $f(A,R_{i})$. This effect is amplified if the region around the optical centre is scarcely populated after $5\sigma$ S/N and colour cuts and suppressed if there are many cluster member candidates (i.e. sources of similar colour). When looking at Fig. 4, it is evident that this effect does not play a role for most of our significant detections. Similarly, this can also happen when using the X-ray centre proxy, if the X-ray centre is misaligned with the optical cluster centre, i.e. the sources closest to the X-ray centre are not necessarily cluster member galaxies. This can be observed when comparing the results using optical vs X-ray centre proxy for 1eRASS JJ140945.2-130101 in Fig. 4 where the two centre proxies are shifted by $1.2\text{\,}\mathrm{a}\mathrm{r}\mathrm{c}\mathrm{m}\mathrm{i}\mathrm{n}$ in RA (compare with Table 1).
@@ 6 Conclusion
In this work we conducted targeted NIR follow-up observations of 17 high-redshift galaxy cluster candidate from eRASS1 and the optically selected high-redshift cluster MOO JJ0319-0025 at the Calar Alto Observatory. We supplemented the target list with additional data of 4 more targets taken with the Wendelstein Wide Field Imager, of which we also observed one at Calar Alto (see Appendix B), totalling to 20 different targets analysed in this work. We performed the prevalent reduction steps using the end-to-end pipeline THELI for the Calar Alto observations and the pipeline by Obermeier et al. (2020) for the Wendelstein observations, and calibrated the $J$-band magnitudes against 2MASS. We complemented the acquired NIR data with the griz-bands of the LSDR10, performed a forced-photometry analysis and detected radial colour over-densities around both X-ray and optical cluster centres in iJi-J, zJz-J, and izi-z for a subsample of our selected targets. For these targets we were able to estimate a photometric redshift. For $2$ of the $20$ different targets no $i$-band data is available from the Legacy Survey at this point in time, thus they were solely analysed in zJz-J. We only obtain two significant detections in zJz-J, therefore our main analysis is based on the iJi-J colour. Out of the $18$ cluster candidates with $i$-band data, we can confirm $12$ targets. If we combine the colour indices izi-z and iJi-J, we can confirm 14 out of 18 targets which have an extended likelihood parameter of $\mathrm{EXT\_LIKE}>5$. This would correspond to a $22.2\%$ contaminant fraction at the high-redshift end of the eRASS1 galaxy cluster catalogue. If we combine the visual inspection classes good and tentative to a confirmed fraction, we would obtain a lenient estimate for the contamination fraction of $25\%$. This would give support to our choices in the used cuts in detections significance ($\mathrm{SIGN}>3\sigma$).
@@
Appendix A Mock Validation
@@ Appendix B CAHA vs WST
Appendix C Supplementary Data Section

### v2

# 3INAF, Osservatorio di Astrofisica e Scienza dello Spazio, via Piero Gobetti 93/3, I-40129 Bologna, Italy

@@
3INAF, Osservatorio di Astrofisica e Scienza dello Spazio, via Piero Gobetti 93/3, I-40129 Bologna, Italy
@@
N. Zimmermann 11    M. Kluge 22    S. Grandis 11    T. Schrabback 11    F. Balzer 22    E. Bulbul 22    J. Comparat 22    B. Csizi 11    V. Ghirardini 33    H. Jansen 11    F. Kleinebreil 11    A. Liu 22    A. Merloni 22    M. E. Ramos-Ceja 22    J. Sanders 22    X. Zhang 22    P. Aschenbrenner 11    F. Enescu 11    S. Keiler 11    M. Märk 11    M. Rinner 11    P. Schweitzer 11    E. Silvestre-Rosello 11    L. Stepman 11
@@ 3.1 Radial colour distribution
where $\mathcal{L_{\mathrm{tot,none}}}$ is the total log-likelihood assuming no additional cluster component. Here, we follow the formalism for a one dimensional Gaussian likelihood, where $\ln{\mathcal{L_{\mathrm{Gauss,1D}}}}=-\frac{1}{2}\left(\frac{\Delta x}{\sigma}\right)^{2}$ and $\ln{\mathcal{L_{\mathrm{Gauss,1D}}^{\mathrm{ref}}}}=\ln{\mathcal{L_{\mathrm{Gauss,1D}}}}\left(\Delta x=0\right)$. Thus we define this quantity in terms of $\sigma$. Only sources with a detection significance $\mathrm{SIGN}\geq 3.0\sigma$ are treated as significant detections. In the case of a clear peak in the posterior distribution, we estimate the uncertainty for each parameter by taking the symmetric $68\%$ interval in parameter space around the retrieved maximum log-likelihood value. If $\mathrm{SIGN}<3\sigma$, we take the $95\%$-percentile of the retrieved posterior distribution as an upper limit on the respective parameter. If the peak of the posterior distribution is within $1\sigma$ of the prior boundary, we report the $5\%$-percentile as a lower boundary (if within $1\sigma$ of the upper prior boundary), and the $95\%$-percentile as an upper boundary (if within $1\sigma$ of the lower prior boundary). We validate the inference framework on mock catalogues, discussed in more detail in Appendix B
@@ 4 Results
In this section, we present the results of the method presented in Sect. 3.1 and the photometric redshift estimation described in Sect. 3.2. Additionally, we provide comments on the visual inspection of the targets analysed in this work. We show a summary of the significances found using the different colour indicies in Fig. 4, using both the X-ray centres and the optical centres provided by Kluge et al. (2024). In addition, Table 3 provides detailed results for the optical centres and iJi-J colours. The findings for the other combinations of cluster centre and colour are provided in Appendix D. Listed in the tables are the maximum posterior values or the upper/lower limit values ($95\%$ / $5\%$-percentile of the posterior distribution) of each of the parameters of our model, the significance of our detection as described in Sect. 3.1, a photometric redshift estimate obtained in this work ($z_{\mathrm{phot},i-J}$, iJi-J indicating the used colour index), the number of sources found to have a cluster member probability higher than $80\%$ ($n_{\mathrm{clm}}^{0.8}$), as well as the best redshift estimate computed by Kluge et al. (2024).
@@ 4.1 Radial colour over-density
We compare the estimated cluster member probabilities from the CAHA data and the the WST data in Fig. 8. The most probable cluster members (upper right corner of Fig. 8) are identified with both of the data sets, whereas the estimated cluster member probability of less probable cluster member galaxies scatters symmetrically around the one-to-one relation. For an image comparison, please see Fig 12 in Appendix C.
@@ 4.3 Visual inspection
For illustration we provide image cutouts around the respective centres (see Appendix D), and comment on each cluster candidate, as well as our respective findings. We order the targets in three different classes defined by a visual assessment and their statistical significance emerging from our analysis:
@@ 5.3 Cluster centre proxies
We test the eRASS1 X-ray centres as reported in Bulbul et al. (2024) (see Appendix D), as well as optical centres found in Kluge et al. (2024) (see Table 3 and Appendix D). In principle, the optical centre is a more suitable centre proxy for our purposes because most of the times it is located at the position of a high probability, bright cluster member galaxy found in earlier works (Kluge et al. 2024). The colour of the central galaxy has an influence of the extracted mean colour, due to the designed radial weighting of the sources’ colours with the weighting function $f(A,R_{i})$ in Eq. (2). Thus, if the cluster centre is placed on an actual cluster member galaxy, the measured colour is weighted accurately by $f(A,R_{i})$ to find the correct red-sequence colour. If the cluster centre is wrongly placed on a random field galaxy, the measured colour does not correspond to the correct red-sequence colour and is mistakenly weighted strongly by $f(A,R_{i})$. This effect is amplified if the region around the optical centre is scarcely populated after $5\sigma$ S/N and colour cuts and suppressed if there are many cluster member candidates (i.e. sources of similar colour). When looking at Fig. 4, it is evident that this effect does not play a role for most of our significant detections. Similarly, this can also happen when using the X-ray centre proxy, if the X-ray centre is misaligned with the optical cluster centre, i.e. the sources closest to the X-ray centre are not necessarily cluster member galaxies. This can be observed when comparing the results using optical vs X-ray centre proxy for 1eRASS JJ140945.2-130101 in Fig. 4 where the two centre proxies are shifted by $1.2\text{\,}\mathrm{a}\mathrm{r}\mathrm{c}\mathrm{m}\mathrm{i}\mathrm{n}$ in RA (compare with Table 1).
@@ 6 Conclusion
In this work we conducted targeted NIR follow-up observations of 17 high-redshift galaxy cluster candidate from eRASS1 and the optically selected high-redshift cluster MOO JJ0319-0025 at the Calar Alto Observatory. We supplemented the target list with additional data of 4 more targets taken with the Wendelstein Wide Field Imager, of which we also observed one at Calar Alto (see Appendix C), totalling to 20 different targets analysed in this work. We performed the prevalent reduction steps using the end-to-end pipeline THELI for the Calar Alto observations and the pipeline by Obermeier et al. (2020) for the Wendelstein observations, and calibrated the $J$-band magnitudes against 2MASS. We complemented the acquired NIR data with the griz-bands of the LSDR10, performed a forced-photometry analysis and detected radial colour over-densities around both X-ray and optical cluster centres in iJi-J, zJz-J, and izi-z for a subsample of our selected targets. For these targets we were able to estimate a photometric redshift. For $2$ of the $20$ different targets no $i$-band data is available from the Legacy Survey at this point in time, thus they were solely analysed in zJz-J. We only obtain two significant detections in zJz-J, therefore our main analysis is based on the iJi-J colour. Out of the $18$ cluster candidates with $i$-band data, we can confirm $12$ targets. If we combine the colour indices izi-z and iJi-J, we can confirm 14 out of 18 targets which have an extended likelihood parameter of $\mathrm{EXT\_LIKE}>5$. This would correspond to a $22.2\%$ contaminant fraction at the high-redshift end of the eRASS1 galaxy cluster catalogue. If we combine the visual inspection classes good and tentative to a confirmed fraction, we would obtain a lenient estimate for the contamination fraction of $25\%$. This would give support to our choices in the used cuts in detections significance ($\mathrm{SIGN}>3\sigma$).
@@
Appendix A Acknowledgements
We would like to thank the whole staff of the Calar Alto Observatory for their support while operating the telescope and making this experience possible. Special thanks to Ana Guijarro, Gilles Bergond, and Alba Fernández-Martín for the quick and informative exchanges we had also after our visits at the site.
We thank Dr. Lindsey Bleem for her contributions to the preparation of the observing proposal.
VG acknowledges the financial contribution from the contracts Prin-MUR 2022 supported by Next Generation EU (M4.C2.1.1, n.20227RNLY3 The concordance cosmological model: stress-tests with galaxy clusters).
The Innsbruck authors acknowledge support provided by the Austrian Research Promotion Agency (FFG) and the Federal Ministry of the Republic of Austria for Climate Action, Environment, Mobility, Innovation and Technology (BMK) via the Austrian Space Applications Programme with grant numbers 899537, 900565, and 911971. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 101004719. This material reflects only the authors views and the Commission is not liable for any use that may be made of the information contained therein. This research has made use of the Spanish Virtual Observatory (http://svo.cab.inta-csic.es) supported by the MINECO/FEDER through grant AyA2017-84089.7 The Wendelstein 2m telescope project was funded by the Bavarian government and by the German Federal government through a common funding process. Part of the 2m instrumentation including some of the upgrades for the infrastructure were funded by the Cluster of Excellence “Origin of the Universe” of the German Science foundation DFG. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. The eROSITA data shown here were processed using the eSASS software system developed by the German eROSITA consortium. E. Bulbul, A. Liu, and X. Zhang acknowledge financial support from the European Research Council (ERC) Consolidator Grant under the European Union’s Horizon 2020 research and innovation program (grant agreement CoG DarkQuest No 101002585). VG acknowledges the financial contribution from the contracts Prin-MUR 2022 supported by Next Generation EU (M4.C2.1.1, n.20227RNLY3 The concordance cosmological model: stress-tests with galaxy clusters). The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID #2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID #2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID #2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory (LBNL). The Legacy Surveys project is honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas A&M University, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientifico e Tecnologico and the Ministerio da Ciencia, Tecnologia e Inovacao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energeticas, Medioambientales y Tecnologicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenossische Technische Hochschule (ETH) Zurich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciencies de l’Espai (IEEC/CSIC), the Institut de Fisica d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig Maximilians Universitat Munchen and the associated Excellence Cluster Universe, the University of Michigan, NSF’s NOIRLab, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas A&M University. BASS is a key project of the Telescope Access Program (TAP), which has been funded by the Nationaal Astronomical Observatories of China, the Chinese Academy of Sciences (the Strategic Priority Research Program “The Emergence of Cosmological Structures” Grant # XDB09000000), and the Special Fund for Astronomy from the Ministry of Finance. The BASS is also supported by the External Cooperation Program of Chinese Academy of Sciences (Grant # 114A11KYSB20160057), and Chinese National Natural Science Foundation (Grant # 12120101003, # 11433005). The Legacy Survey team makes use of data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), which is a project of the Jet Propulsion Laboratory/California Institute of Technology. NEOWISE is funded by the National Aeronautics and Space Administration. The Legacy Surveys imaging of the DESI footprint is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH1123, by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract; and by the U.S. National Science Foundation, Division of Astronomical Sciences under Contract No. AST-0950945 to NOAO.
Appendix B Mock Validation
@@ Appendix C CAHA vs WST
Appendix D Supplementary Data Section